OpenStudy (anonymous):

factor please!!!!!!!!! 5x^2-22x+8

OpenStudy (anonymous):

\[5X^2-22x+8\]

OpenStudy (anonymous):

To factor this, we cannot simplify by distribution so. Multiply a * c. 5*8=40 Now we need two numbers that have a product of 40 and a sum of -22. -20*-2=40 -20+-2=-22 Replace -20x-2x for -22x. Separate by grouping. (5x^2-20x)(-40x+8) We can simplify these factors by factoring out some numbers. 5x(x-4)*-8(x-4) *Note that they both have (x+4) in common* Combine 5x and -8 since they have a common factor. (5x-8)(x-4) Aaand You're done. good luck (:

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome.